# Data driven discovery of human mobility models

### v1

@@ Abstract
Human mobility is a fundamental aspect of social behavior, with broad applications in transportation, urban planning, and epidemic modeling. However, for decades new mathematical formulas to model mobility phenomena have been scarce and usually discovered by analogy to physical processes, such as the gravity model and the radiation model. These sporadic discoveries are often thought to rely on intuition and luck in fitting empirical data. Here, we propose a systematic approach that leverages symbolic regression to automatically discover interpretable models from human mobility data. Our approach finds several well-known formulas, such as the distance decay effect and classical gravity models, as well as previously unknown ones, such as an exponential-power-law decay that can be explained by the maximum entropy principle. By relaxing the constraints on the complexity of model expressions, we further show how key variables of human mobility are progressively incorporated into the model, making this framework a powerful tool for revealing the underlying mathematical structures of complex social phenomena directly from observational data.
{affiliations}
Institute of Remote Sensing and Geographical Information Systems, School of Earth and Space Sciences, Peking University, Beijing, China
These authors contributed equally: Hao Guo, Weiyu Zhang.
Email:leidong@pku.edu.cn; liuyu@urban.pku.edu.cn
@@
A deeper understanding of population movement and its underlying mechanisms is crucial for informed urban planning[1], transportation system management[2], and epidemic control strategies[3, 4]. Historically, the study of human mobility has evolved from qualitative description of key factors [5] to quantitative analyses, such as the characterization of the distance decay effect[6], culminating in formal mathematical models of population flows. The most prominent of these is the gravity model, proposed in the early 20th century[7, 8]. Analogous to Newton’s law of gravitation, the gravity model posits that human mobility flows are proportional to the population of interacting locations and inversely related to the distance between them[9, 10].
@@
Despite its enduring influence, progress in developing analytical models for population flows has been intermittent. Notable contributions include the intervening opportunity model[11, 12], the radiation model[13], and models incorporating preferential return[14, 15, 16]. However, a major obstacle to further progress within this physically-informed paradigm has been the lack of analytical frameworks derived from first principles. Unlike the physical sciences where models are often grounded in fundamental laws, social science models — particularly those describing mobility — are predominantly empirical, relying on intuition and interactive exploration to identify appropriate model forms. The inherent complexity and heterogeneity of human movement further compound this challenge.
@@
The recent availability of high-resolution mobility data from sources such as call detail records (CDR), real-time smartphone tracking, and social media check-ins[17, 18, 19] has offered unprecedented opportunities to investigate human mobility. These datasets provide valuable insights into both routine commuting patterns and more sporadic travel behaviors. However, the high dimensionality and inherent noise in these datasets pose substantial challenges for model identification using traditional, intuition-based approaches[20]. This necessitates automated methods capable of leveraging large observational datasets to accelerate the discovery of accurate and generalizable models of human mobility.
@@
Here, we introduce a machine learning-based approach for discovering models of human mobility. Specifically, we use Symbolic Regression (SR)[21], a powerful technique that can uncover analytical expressions directly from data [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Unlike traditional regression methods, SR does not assume a pre-specified function forms. Instead, it learns both the function structure and the associated parameters[32]. This flexibility enables the discovery of novel scientific laws[33] and provides insights into the underlying mechanisms of complex phenomena. Moreover, SR has shown the ability to produce more generalizable models than other machine learning techniques, such as neural networks[34, 35], due to its inherent preference for simpler, more compact expressions.
@@
We apply this approach to mobility data from three countries - China, the UK, and the US - examining both commuting and general mobility flows. Our results show that the exponential decay gravity model consistently emerges across countries and spatial scales, suggesting its optimality in describing human mobility under certain model complexities (i.e. expression length). Beyond recovering known models, SR uncovers novel extensions, including a previously unknown exponential power-law distance decay, which can be interpreted through the maximum entropy principle[36]. Furthermore, by analyzing symbolic models derived from flow data grouped by geographic regions, we reveal notable heterogeneity in model forms between intra- and inter-region flows, providing new insights into the diverse modes of human movement that are often masked in global models. Finally, using synthetic flow data, we validate the robustness of SR to discover various mobility models from noisy observations.
Fig. 1: The analytical framework of mobility model distillation. (a) Mobility flow of Guangdong, China. The flow volume $F_{ij}$ from origin $i$ to destination $j$ is the response variable. (b) The explanatory variables include the workplace population $w$, the residential population $r$, geographic distance $d_{ij}$, and intervening opportunities $s_{w},s_{r}$, calculated with workplace and residential population, respectively. (c) The overall workflow to automatically distill models from mobility data. Starting with response variables and explanatory variables along with seven common operators, we use a genetic-programming-based SR program to search for appropriate model forms. In each iteration, the SR program generates models for allocation weight function $f$. For each origin $i$, the total outflow $O_{i}$ is allocated to each destination based on the corresponding allocation weights $f_{ij}$. The MSE between the allocated flow $\hat{F}_{ij}$ and the actual flow $F_{ij}$ is then calculated and fed back into the SR program for expression optimization.
Results
0.1 Research design
@@
To distill mathematical models of human mobility, we first compiled four large-scale datasets from three countries: 2020 cellphone data from Guangdong Province, China; 2019 cellphone data from Beijing-Tianjin-Hebei urban agglomeration, China; 2011 Census data from England; and 2011-2015 American Community Survey (ACS) data from the US (see Methods for details). These datasets encompass commuting flows (both cellphone and survey data) and general mobility patterns (cellphone data) across various spatial resolutions and geographical areas, thereby enabling a comprehensive evaluation of the robustness of potential mobility models.
@@
Figure 1 presents the analytical framework. For each dataset, we consider four types of explanatory variables: workplace population $w$, residential population $r$, geographical distance $d$, and intervening opportunities $s$ (Fig. 1b). Here, intervening opportunities represent the number of competing destinations that are closer to the origin[13, 17]. In line with the literature, we use the sum of the population in these areas as a proxy[13]. Besides variables, we consider five commonly used binary operators (Fig. 1c): addition ($+$), subtraction (--), multiplication ($\times$), division ($\div$), power (^); and two unary operators: exponentiation ($\exp$), logarithm ($\ln$). The goal of the SR algorithm is to identify the optimal model form by simultaneously optimizing the accuracy and complexity. Specifically, accuracy is measured by the Mean Squared Error (MSE) of the flow volumes over the entire dataset, while complexity is quantified by the length of the expression, defined as the total number of variables, operators, and constants, with repeated symbols counted separately [20]. Each constant, variable, and binary operator is assigned a length of 1, while $\exp$ and $\ln$ are assigned a length of 2 due to the implicit presence of the constant $e$. When the expression is represented as a full binary tree, model complexity corresponds to the number of nodes, which must always be an odd integer (see Extended Data Fig. 1 for examples).
Rather than directly modeling the origin-destination flow $F_{ij}$ between places $i$ and $j$, we model the allocation weight $f_{ij}$ with SR. This weight is proportional to the probability that an individual starting from place $i$ will choose destination $j$ (Fig.  1c). Let $O_{i}$ denote the total outflow from place $i$, $F_{ij}$ is calculated from $\{f_{ik}\}_{k\neq i}$: $\hat{F}_{ij}=O_{i}f_{ij}/\sum_{k}f_{ik}$. This approach reduces the length of the target expression. For example, the power-law decay gravity model $F_{ij}=km_{i}m_{j}/d_{ij}^{\beta}$ has a complexity of 9, while its allocation weight form $f_{ij}=m_{j}/d_{ij}^{\beta}$ has a complexity of 5, where $m_{i}$, $m_{j}$ are the populations of the origin $i$ and destination $j$, respectively. The search space of SR grows exponentially with expression length, which can largely impact its performance on complex models. Therefore, the allocation weight approach enhances the stability and efficiency of SR.
@@
To identify the optimal model form, we use a Genetic Programming-based SR algorithm (Fig. 1c). This algorithm evolves a population of expressions represented as trees. In analogy to natural selection, expressions with higher fitness values (calculated based on accuracy and complexity) have a higher probability of producing offspring through crossover and mutation. The algorithm generates a Pareto frontier of optimal expressions, with each expression being Pareto-optimal if no improvement in one objective (accuracy or complexity) can be made without sacrificing the other. Detailed SR settings are provided in the Methods section.
Fig. 2: SR results on mobility flow data. (a-c) Pareto frontiers of SR models on Guangdong, England, and US datasets. As flow magnitudes vary across datasets, we normalize the RMSE with that of the simplest gravity model ($m_{j}/d_{ij}$). The accuracy and complexity of six existing models are marked with crosses (note that some existing models are not shown as their errors exceed the range of the y-axis). Expressions with complexity 1 are excluded as one variable or constant is not sufficient to model the mobility flow. Across all datasets and complexity levels, the SR models consistently outperform or match the existing models in accuracy. (d) Expressions of the Pareto optimal SR models. The notations follow Fig. 1, except that $d$ is short for $d_{ij}$ and $w,r$ for $w_{j},r_{j}$ (as origin populations are not present in these formulas). To align with existing models, some expressions are not given in the form with the lowest complexity. The distilled models are classified based on the captured effects governing human mobility. (e) Expressions of the considered existing models. GMZipf: Zipf’s simple gravity model[8]; GMPow/GMExp: gravity model with power-law/exponential decay[36]; OPS: opportunity priority selection model[37]; RM: radiation model[13]; IO: Schneider’s intervening opportunity model[12].

### v2

# Distilling human mobility models with symbolic regression

@@ Abstract
Human mobility is a fundamental aspect of social behavior, with broad applications in transportation, urban planning, and epidemic modeling. Represented by the gravity model and the radiation model, established analytical models for mobility phenomena are often discovered by analogy to physical processes. Such discoveries can be challenging and rely on intuition, while the potential of emerging social observation data in model discovery is largely unexploited. Here, we propose a systematic approach that leverages symbolic regression to automatically discover interpretable models from human mobility data. Our approach finds several well-known formulas, such as the distance decay effect and classical gravity models, as well as previously unknown ones, such as an exponential-power-law decay that can be explained by the maximum entropy principle. By relaxing the constraints on the complexity of model expressions, we further show how key variables of human mobility are progressively incorporated into the model, making this framework a powerful tool for revealing the underlying mathematical structures of complex social phenomena directly from observational data.
Keywords: Symbolic Regression, human mobility, spatial interaction, gravity model
1 Introduction
@@
A deeper understanding of population movement and its underlying mechanisms is crucial for informed urban planning [1, 58], transportation system management [8], and epidemic control strategies [15, 38]. Historically, the study of human mobility has evolved from qualitative description of key factors [34] to quantitative analyses, such as the characterization of the distance decay effect [25], culminating in formal mathematical models of population flows. The most prominent of these is the gravity model, proposed in the early 20th century [47, 62]. Analogous to Newton’s law of gravitation, the gravity model posits that human mobility flows are proportional to the population of interacting locations and inversely related to the distance between them [36, 2].
@@
Despite the enduring influence of the class of gravity models [57], progress in developing analytical models for population flows has been intermittent. Notable contributions include the intervening opportunity model [48, 41], the radiation model [45], and models incorporating preferential return [46, 4, 39]. A major challenge within this physics-informed paradigm has been the lack of analytical frameworks derived from first principles. Unlike the physical sciences where models are often grounded in fundamental laws, models describing social phenomena, such as mobility behavior, are predominantly empirical, relying on intuition and interactive exploration to identify appropriate model forms. The inherent complexity and heterogeneity of human movement further compound this challenge.
On the other hand, there has been rapid development of machine-learning models for human mobility, especially those based on artificial neural networks [44, 28] and graph neural networks [29, 60, 24]. These models typically incorporate multiple features (e.g., points of interest, road density, land use) to achieve better prediction accuracy. Yet their contribution to understanding human mobility has been limited due to their black-box nature.
@@
The recent availability of high-resolution mobility data from sources such as call detail records (CDR), real-time smartphone tracking, and social media check-ins [3, 55, 32] has offered unprecedented opportunities to investigate human mobility. These datasets provide valuable insights into both routine commuting patterns and more sporadic travel behaviors. However, the high dimensionality and inherent noise in these datasets pose substantial challenges for model identification using traditional, intuition-based approaches [9]. This necessitates automated methods capable of leveraging large observational datasets to accelerate the discovery of accurate and interpretable models of human mobility.
@@
The endeavor to build symbolic mobility models automatically dates back to [31] and [11]. Nevertheless, the potential of this paradigm was not fully realized due to inefficient mobility data and limitations in model optimization methods at that time. Without proper regularization of model complexity, the discovered expressions are lengthy and difficult to interpret. Nowadays, Symbolic Regression (SR) has emerged as a powerful technique to uncover analytical expressions directly from data [30], with applications in diverse scientific disciplines [35, 56, 54, 24, 51, 27]. SR learns both the function structure and the associated parameters, relieving the need to specify the function form a priori [20]. This flexibility enables the discovery of novel scientific laws and provides insights into the underlying mechanisms of complex phenomena [7]. Moreover, SR has shown the ability to produce more generalizable models than other machine learning techniques, such as neural networks, due to its inherent preference for simpler, more compact expressions [52, 21, 6].
@@
Here, we propose a Symbolic Regression-based framework for discovering models of human mobility. We apply this approach to mobility data from three countries - China, the UK, and the US - examining both commuting and general mobility flows. Our results show that the exponential decay gravity model consistently emerges across countries and spatial scales, suggesting its optimality in describing human mobility under certain model complexities (i.e. expression length). Beyond recovering known models, SR uncovers novel extensions, including a previously unknown exponential power-law distance decay, which can be interpreted through the maximum entropy principle [57]. Furthermore, by analyzing symbolic models derived from flow data grouped by geographic regions, we reveal notable heterogeneity in model forms between intra- and inter-region flows, providing new insights into the diverse modes of human movement that are often masked in global models. Finally, using synthetic flow data, we validate the robustness of SR to discover various mobility models from noisy observations.
2 Research design
2.1 The analytical framework
@@
Figure 1 presents the analytical framework. To distill mathematical models of human mobility, we first compiled four large-scale datasets from three countries: 2020 cellphone data from Guangdong Province, China; 2019 cellphone data from Beijing-Tianjin-Hebei urban agglomeration, China; 2011 Census data from England; and 2011-2015 American Community Survey (ACS) data from the US (see Section 2.2 for details). These datasets encompass commuting flows (both cellphone and survey data) and general mobility patterns (cellphone data) across various spatial resolutions and geographical areas, thereby enabling a comprehensive evaluation of the robustness of potential mobility models.
Figure 1: The analytical framework of mobility model distillation. (a) Mobility flows in Guangdong, China. The flow volume $F_{ij}$ from origin $i$ to destination $j$ is the response variable. (b) The explanatory variables include the workplace population $w_{i},w_{j}$, the residential population $r_{i},r_{j}$, geographic distance $d_{ij}$, and intervening opportunities $s^{(w)}_{ij},s^{(r)}_{ij}$, calculated with workplace and residential population, respectively. (c) The overall workflow to automatically distill models from mobility data. Considering seven common operators, we use a genetic-programming-based SR program to search for appropriate model forms. In each iteration, the SR program generates models for allocation weight function $f$. For each origin $i$, the total outflow $O_{i}$ is allocated to each destination based on the corresponding allocation weights $f_{ij}$. The MSE between the predicted flow $\hat{F}_{ij}$ and the actual flow $F_{ij}$ is then calculated and fed back into the SR program for expression optimization.
Given a dataset with spatial units indexed by $i=1,2,\dots,n$, we consider four types of explanatory variables: workplace population $w_{i}$, residential population $r_{i}$, geographical distance $d_{ij}$, and intervening opportunities $s_{ij}$ (Fig. 1b). Here, intervening opportunities represent the number of competing destinations that are closer to the origin [3]. In line with the literature, we use population as a proxy of competing destinations [45]. As we consider both residential and workplace population, we calculate intervening opportunities based on each of them:
$s_{ij}^{(r)}$ $=\sum_{k:0<d_{ik}<d_{ij}}r_{k}$ (1)
$s_{ij}^{(w)}$ $=\sum_{k:0<d_{ik}<d_{ij}}w_{k}.$ (2)
Besides variables, we consider five basic binary operators (Fig. 1c): addition (++), subtraction (-), multiplication ($\times$), division ($\div$), power (ˆ); and two unary operators: exponentiation ($\exp$), logarithm ($\ln$).
@@
The goal of symbolic regression is to identify the optimal model form by simultaneously optimizing the accuracy and complexity. Specifically, accuracy is measured by the mean squared error (MSE) of the flow volumes over the entire dataset, while complexity is quantified by the length of the expression, defined as the total number of variables, operators, and constants, with repeated symbols counted separately [9]. Each constant, variable, and binary operator is assigned a length of 1, while $\exp$ and $\ln$ are assigned a length of 2 due to the implicit presence of the constant $e$. When the expression is represented as a full binary tree, model complexity corresponds to the number of nodes, which must always be an odd integer (see Fig. 2 for examples).
Figure 2: Examples for calculating the model complexity with binary expression trees. (a) Under the allocation weight setting (see Section 2.3), the gravity model with power-law decay $f_{ij}=m_{j}/d_{ij}^{\beta}$ is expressed as a tree with 5 nodes, hence has a complexity of 5. (b) The radiation model $f_{ij}=m_{j}/(m_{i}+s_{ij})(m_{i}+s_{ij}+m_{j})$ has a complexity of 11.
@@
To identify the optimal model form, we apply a genetic programming-based SR algorithm (Fig. 1c). This algorithm evolves a population of expressions represented as trees. In analogy to natural selection, expressions with higher fitness values (calculated based on accuracy and complexity) have a higher probability of producing offspring through crossover and mutation. The algorithm generates a Pareto frontier of optimal expressions, with each expression being Pareto-optimal if no improvement in one objective (accuracy or complexity) can be made without sacrificing the other. Detailed SR settings are given in Section 2.3.
We use the root mean squared error (RMSE) and common part of commuters (CPC) [23] to measure the discrepancy between the model-predicted flows and real flows. CPC is a common metric for models of origin-destination flows [22, 44]. It is defined as:
$\text{CPC}=\frac{2\sum_{i,j}\min\{F_{ij},\hat{F}_{ij}\}}{\sum_{i,j}F_{ij}+\sum_{i,j}\hat{F}_{ij}},$ (3)
where $F_{ij},\hat{F}_{ij}$ are the real and predicted flow volume from place $i$ to $j$, respectively. The range of CPC is $[0,1]$, and higher value indicates a better match between real and predicted flows.
We compare the SR-derived models with six existing mobility models (Fig.3e): Zipf’s simple gravity model (GMZipf, [62]); gravity model with power-law/exponential decay (GMPow/GMExp, [57]); opportunity priority selection model (OPS, [26]); radiation model (RM, [45]); Schneider’s intervening opportunity model (IO, [41]). These models are also calibrated to minimize MSE. For commuting datasets, we use workplace population (and intervening opportunities based on workplace population) in the existing models, as using residential population yields inferior results.
2.2 Data
A general description of the flow datasets is given in Table 1. The mobility flows are visualized in Supplementary Figs. S2-S6.
The Guangdong dataset is provided by China Mobile Communications Corporation, the largest telecommunications operator in China. The raw dataset comprises individual movement trajectories of 5 million users in November 2020, with a spatial resolution of 500 m $\times$ 500 m. After filtering out incorrect or duplicate trajectory points, we identify stay points as trajectory points with a duration of at least 30 min. The mobility flows are movements between stay points, aggregated at the sub-district and county levels based on the grid cells’ center locations.
For the commuting flow dataset, an individual’s home is defined as the location where he stayed the longest between 12:00 AM and 7:00 AM, and the workplace is the location where he stayed the longest between 9:00 AM and 6:00 PM. Commuting trips are identified as movements from the home location to the workplace in the morning (before 12:00 PM). Subsequently, the commuting movements are aggregated similarly to the mobility flow dataset. The residential / workplace populations are derived based on the detected home and workplace locations.
The Beijing-Tianjin-Hebei mobility data is provided by China Unicom, which contains inter-county mobility from November 4 to 10, 2019. Similarly, the mobility flows are aggregated movements between stay points with durations $\geq$ 30 min. Registered population data in 2019 are collected from official statistical yearbooks.
The commuting datasets in England and the Contiguous US record the location of usual residence and workplace of employed residents, aggregated to merged local authority district (MLAD) and county levels, respectively. The sources of residential and workplace population data are the same as the flow data (2011 Census for England, and 2011-2015 ACS 5-year estimate for US).
For all datasets above, we only consider flows between spatial units, so flows with the same origin and destination are removed. Moreover, we use the population data at finer scale (if available) to calculate the intervening opportunity. For example, to calculate the intervening opportunities from county $i$ to $j$ in Guangdong, we add up the population of all sub-districts $k$ outside $i$ and $j$ where $d_{ik}<d_{ij}$. Similarly, population at middle-layer super output area (MSOA) level is used for the England dataset.
Table 1: Description of the real flow datasets. Area and population are the median values among all spatial units. An observation refers to an origin-destination pair. BTH is short for Beijing-Tianjin-Hebei urban agglomeration, China.
Region Granularity Type #Units #Observations Area ($\text{km}^{2}$) Population
Guangdong sub-district mobility 2,100 705,811 60 36,962
commuting 28,392
Guangdong county mobility 124 14,251 1,256 725,391
commuting 1,150
England MLAD commuting 324 57,591 212 126,430
US county commuting 3,108 131,391 1,616 25,991
BTH county mobility 200 38,813 756 440,000
2.3 Symbolic Regression
Rather than directly modeling the origin-destination flow $F_{ij}$ between places $i$ and $j$, we model the allocation weight $f_{ij}$ with symbolic regression. This weight is proportional to the probability that an individual starting from place $i$ will choose destination $j$ (Fig. 1c). Let $O_{i}$ denote the total outflow from place $i$, $F_{ij}$ is calculated from $\{f_{ik}\}_{k\neq i}$ as follows:
$\hat{F}_{ij}=O_{i}\frac{f_{ij}}{\sum_{k\neq i}f_{ik}}.$ (4)
This approach reduces the complexity of the target models. For example, the power-law decay gravity model
$F_{ij}=km_{i}m_{j}/d_{ij}^{\beta}$ (5)
has a complexity of 9, while its allocation weight form
$f_{ij}=m_{j}/d_{ij}^{\beta}$ (6)
has a complexity of 5, where $m_{i}$, $m_{j}$ are the populations of the origin $i$ and destination $j$, respectively. The search space of SR grows exponentially with expression length, which can largely impact the performance of SR to find the optimal function form. Our allocation weight approach reduces the computation burden of SR, enhancing its stability and efficiency.
Based on the allocation weight approach, our SR problem can be formulated as a minimisation problem to find
$f^{*}\in\mathop{\arg\min}\limits_{f\in\mathcal{F}}\frac{1}{n(n-1)}\sum_{i,j}\left(F_{ij}-O_{i}\frac{f\left(r_{i},r_{j},w_{i},w_{j},d_{ij},s^{(r)}_{ij},s^{(w)}_{ij}\right)}{\sum_{k\neq i}f\left(r_{i},r_{k},w_{i},w_{k},d_{ik},s^{(r)}_{ik},s^{(w)}_{ik}\right)}\right)^{2}+\lambda C(f),$ (7)
where $\mathcal{F}$ is the set of functions which can be formed by composition of the allowed operators; $n$ is the number of spatial units; $C(f)$ is the complexity of the function $f$; $\lambda$ is a penalty factor for model complexity. Such a trade-off between model accuracy and simplicity is typical in statistical model selection [5]. Furthermore, if we formulate SR as a multi-objective optimization problem regarding accuracy and simplicity, changing the parameter $\lambda$ would lead to different models on the Pareto frontier.
SR is proven to be NP-hard [53], which indicates an exact and efficient solution probably does not exist. As a result, the heuristic genetic programming approach, [18, 40, 20] remains the mainstream solution to SR, though methods based on physical principles [50, 49], neural networks [37, 33, 17], and Bayesian computation [14, 16] are emerging recently.
We develop our SR program based on a Julia package for genetic programming based SR, SymbolicRegression.jl [9], as it is specially designed for scientific data analysis. Instead of penalizing the complexity of each expression (Eq.7), this software applies a penalty on complexity levels if they are too frequent in the population. In other words, it aims to balance explorations for models at a range of complexity (from 1 to 19 in this study). As a result, it can output the whole Pareto frontier in a single run. Moreover, expression simplification and constant optimization are included in the set of mutation operations to improve the models’ compactness and accuracy. Several other enhancements to the general genetic programming framework are implemented, including the integration of simulated annealing, age-regularization to avoid premature convergence, and migration of expressions between several parallel populations. We refer to [9] for more technical details.
Nevertheless, optimization of the allocation weight function is not readily supported by the package. We modified the evaluation module in SymbolicRegression.jl to calculate flow volumes from allocation weights. As constant optimization only occurs with probability, we also added a constant optimization step before the program exits, ensuring all output formulas have optimized parameter values.
Due to the randomness of genetic programming, we repeat the SR procedure multiple times for each dataset. Typically, we first use all features as explanatory variables, then restrict the set of variables to those shown to be important. Fewer variables lead to a smaller search space, thus may yield better solutions. The number of iterations is set as 100 throughout this study. Batching is necessary to accelerate the program for large flow datasets. By default, the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm is used to optimize constants, and the Nelder-Mead algorithm is used if the former fails. The Pareto optimal expressions within the combined output of repeated runs are selected as final models. The detail of SR settings on each dataset is given in Supplementary Table S1.
3 Results